# Robustness in Large Language Models: A Survey of Mitigation Strategies and Evaluation Metrics

### v1

@@ 4.4 Post-processing techniques
A fundamental principle of post-processing is to treat the LLMs output with a degree of skepticism, applying verification steps similar to how one might validate user input. Common strategies include safety classifiers, where smaller models or rule-based systems screen outputs for harmful or biased content, blocking or flagging problematic responses (Kumar et al., 2025b; Liu et al., 2023; Zhou et al., 2024a). For example, the Erase-and-Check method iteratively tests modified versions of suspicious prompts to detect adversarial inputs, leveraging pre-generation filtering to block harmful content (Kumar et al., 2025a; Phute et al., 2024). Perplexity filtering identifies nonsensical outputs by measuring how natural generated text appears to a language model, flagging high-perplexity results as potential errors or attacks (Weng, 2023).
@@
For structured tasks (e.g., code or JSON generation), format validation ensures outputs adhere to predefined schemas or pass external checks like code compilers, preventing errors or injection vulnerabilities (Rebedea et al., 2023; AI, 2025). Similarly, output sanitization removes sensitive information (e.g., personal data) leaked during generation, mitigating privacy risks (Wang et al., 2025). Further, Song et al. (2025); Maheshwari et al. (2025) implements RAG to verify outputs after initial response generation. After a language model produces an answer, the system automatically fetches supporting evidence from external databases or documents, using either the original user query or the generated response as search criteria. The model then cross-checks its own output against this retrieved evidence, creating a self-correction cycle that improves factual consistency. While methods like Erase-and-Check offer strong defences, their computational cost often requires simplified variants for practical use, highlighting the trade-off between robustness and efficiency. Together, these strategies underscore the importance of layered verification to balance safety, accuracy, and usability in real-world LLM deployments.
@@
This approach uses one language model to evaluate the outputs of another, offering scalable and nuanced assessments beyond basic rule-based checks. This method is increasingly applied to enhance robustness by automating tasks like bias detection, safety compliance, and factual accuracy (Gu et al., 2025; Kumar et al., 2025b). For instance, judge LLMs can systematically analyze responses for fairness across demographic categories, validate reasoning steps in technical tasks (e.g., code vulnerability detection), or flag factual errors by cross-referencing trusted sources (Cantini et al., 2025). Advanced frameworks even deploy judge LLMs to autonomously generate adversarial prompts, test target models, and assess response reliability, guided by domain-specific constraints (Li et al., 2025c; Beyer et al., 2025).
However, reliance on LLM judges introduces challenges. First, their evaluations can be inconsistent due to prompt sensitivity, inherent biases, or a lack of grounding in objective criteria (Gu et al., 2025). Second, judge models themselves are prone to manipulation; studies show adversarial phrases or stylistic cues (e.g., phrases like “I think”) can skew their assessments (Raina et al., 2024; Lee et al., 2024a). Finally, the lack of standardized evaluation protocols and risks like “preference leakage”, where judges favour outputs from models they were trained on, underscore the need for multi-judge ensembles, human oversight, and rigorous benchmarking to ensure reliability (Beyer et al., 2025). While promising, LLM-as-a-Judge systems require careful design to balance automation with trustworthiness.
@@ 5.1 Metrics
Performance Degradation Accuracy Drop, F1 Drop (Yuan et al., 2023) General Robustness Decrease in standard metrics on challenging vs clean data.
Attack Success Rate (ASR) (Zhou et al., 2024a) Adversarial Resilience, Safety Percentage of successful adversarial attacks/jailbreaks
Compliance / Refusal Rate (Varshney et al., 2024; Zeng et al., 2025) Safety, Alignment Frequency of generating harmful content or refusing prompts
OOD Performance OOD Accuracy/F1 score (Yuan et al., 2023) OOD Generalization Task performance on designated out-of-distribution datasets.
ID-OOD Performance Gap (Yuan et al., 2023) OOD Generalization Difference between in-distribution and out-of-distribution performance.
OOD Detection (AUROC, Energy Score) (Lu et al., 2024) OOD Awareness Ability to distinguish in-distribution versus out-of-distribution inputs.
Consistency Semantic Consistency (Yang et al., 2025b) Prompt Robustness, Reliability Stability of output meaning under input paraphrasing or slight modifications.
Consistency Rate (CR) (Nalbandyan et al., 2025) Prompt Robustness, Sampling Stability Proportion of identical or equivalent predictions across setups (e.g., prompts, seeds).
@@
(RC) (Rahman et al., 2024) paraphrased prompts.
@@
Consistency Score (Saxena et al., 2024) Reliability samples.
Calibration Expected Calibration Error (ECE) (Posocco & Bonnefoy, 2021) Confidence Reliability Alignment between predicted confidence and actual accuracy
ECE under Distribution Shift (Huang et al., 2024b; LeVine et al., 2024) OOD Generalization ECE measured specifically on OOD data
Fairness under Stress Bias Metrics (e.g., FLEX) (Jung et al., 2025) Fairness Robustness Evaluation of fairness when models are exposed to bias-inducing prompts.
Task-Specific Code Robustness Checks (Li et al., 2025d) Code Generation Robustness Checks for robustness in generated code, like security or correctness features.
Reasoning Accuracy (e.g., ReClor-plus) (Bao et al., 2025) Reasoning Robustness Performance on reasoning tasks containing structure variations.
@@
(Rahman et al., 2024) Faithfulness knowledge source or context.
FEWL (Wei et al., 2024) Factuality (Reference-Free) Weighted agreement among reference LLMs based on expertise.
SelfCheckGPT, MetaQA (Wei et al., 2024) Factuality/Consistency (Reference-Free) (Chen et al., 2024) Consistency checks across multiple generated samples or metamorphic prompt variations.
@@ 5.1.1 Performance Degradation Metrics
These metrics quantify the drop in performance when an LLM is evaluated on challenging data (adversarial, OOD, noisy) compared to its performance on clean, standard (IID) data. They provide a direct measure of how much a specific challenge impacts the model’s effectiveness.
  • Accuracy Drop / F1 Drop / Metric Drop: These are calculated as the difference (absolute or percentage) between a standard metric (like Accuracy or F1-score) on the IID test set and the same metric on the challenging test set (e.g., an OOD dataset or an adversarially perturbed dataset) (Yuan et al., 2023; Du et al., 2023). A larger drop indicates lower robustness to that specific challenge. For example, a 20% accuracy drop on an OOD dataset suggests poor generalization.
  • Attack Success Rate (ASR): This metric is central to adversarial robustness evaluation. It measures the percentage of adversarial inputs that successfully cause the intended model failure, such as misclassification in NLU tasks, bypassing safety filters to generate harmful content, or successfully executing a jailbreak prompt (Jung et al., 2025). A higher ASR signifies lower robustness to the specific attack method used. It is a key metric for benchmarks like AdvGLUE++, JailbreakBench, and PromptBench (Zhou et al., 2024a).
  • Compliance Ratio / Refusal Rate: These metrics are particularly relevant for evaluating safety and alignment robustness, especially against jailbreaking attempts. It measures how often the LLM generates harmful or forbidden content when prompted with malicious instructions (Varshney et al., 2024; Zeng et al., 2025). Conversely, the Refusal Rate measures how often the model refuses to answer. A high refusal rate for harmful prompts indicates good safety alignment, but a high refusal rate for benign prompts indicates over-defensiveness, a potential negative side-effect of safety training.
@@ 5.1.2 Out-of-Distribution (OOD) Performance Metrics
  • OOD Accuracy / F1 score: Standard performance metrics like Accuracy or F1-score calculated directly on designated OOD benchmark datasets, such as those included in the BOSS suite (e.g., Dynasent, ANLI, NewsQA) (Yuan et al., 2023). This provides a direct measure of capability on unseen distributions.
  • ID-OOD Performance Gap: This metric quantifies the difference in performance between an ID dataset and a corresponding OOD dataset (Yuan et al., 2023). It explicitly measures the drop in generalization capability due to the distribution shift. Studies using benchmarks like BOSS have shown that this gap can vary significantly and exhibit complex patterns (e.g., linear, piecewise linear, V-shaped) depending on factors like model scale, training steps, and task type, indicating that simply improving ID performance does not always guarantee better OOD performance.
  • OOD Detection Metrics (e.g., AUROC, FPR@TPR): These metrics evaluate the model’s ability to recognize that an input is OOD, often by thresholding a confidence score derived from model outputs (e.g., token probabilities) or internal representations (Huang et al., 2024a). Common metrics include the Area Under the Receiver Operating Characteristic curve (AUROC) and the False Positive Rate (FPR) at a high True Positive Rate (TPR). While not directly measuring task performance robustness, effective OOD detection is crucial for building reliable systems that can identify and potentially abstain from making predictions on inputs they are likely to handle poorly (Lu et al., 2024). Techniques like Energy Score, originally from classification, have been adapted for this purpose in LLMs and reward models (LeVine et al., 2024).
@@ 5.1.3 Consistency Metrics
Consistency metrics evaluate the stability of LLM outputs. A consistent model should produce similar outputs for semantically identical inputs and exhibit stable behaviour under non-deterministic sampling. Lack of consistency can signal unreliability, especially for tasks requiring factual accuracy (Nalbandyan et al., 2025).
  • Semantic Consistency: Assesses whether the meaning or core information conveyed in the LLM’s output remains the same when the input prompt is paraphrased or expressed differently, while preserving the original intent (Nalbandyan et al., 2025). Evaluation can be qualitative or quantitative, often using semantic similarity metrics based on embeddings (e.g., cosine similarity between output embeddings) (Patwardhan et al., 2025).
  • Consistency Rate (CR): Proposed within the SCORE framework (Nalbandyan et al., 2025), CR quantifies the agreement among predictions for the same input across different evaluation settings (e.g., 10 different prompts, 5 different random seeds for sampling). It is calculated as the average proportion of agreeing pairs among all possible pairs of predictions for each input instance. For multiple-choice questions (MCQ), agreement means identical predicted letters; for tasks like MATH, it means symbolic equivalence checked via tools like sympy (Nalbandyan et al., 2025).
  • Response Consistency (RC): Used in the DefAn benchmark (Bao et al., 2025), this metric specifically measures consistency across multiple (e.g., 15) paraphrased versions of a single prompt. For each set of paraphrased prompts, it calculates the frequency of the most common factual claim generated in the responses. The overall RC is the average of these frequencies.
  • Multilingual Consistency: Evaluates whether the model’s performance or the semantic meaning of its output is preserved when prompts or inputs are translated into different languages (Qi et al., 2023).
  • LLM-based Consistency Score: This approach uses another LLM as a judge (Saxena et al., 2024). It prompts the target LLM multiple times (n > 1) for the same input using a non-zero temperature (e.g., 1.0) to induce variability. The judge LLM then compares pairs of generated responses, assessing whether they contradict or support each other. The consistency score is the fraction of pairs deemed consistent.
@@ 5.1.4 Calibration Metrics
  • Expected Calibration Error (ECE): It works by dividing predictions into bins based on their confidence scores (e.g., 10 bins from 0.0-0.1, 0.1-0.2,…, 0.9-1.0). Within each bin, the average confidence and the actual accuracy (fraction of correct predictions) are calculated. ECE is the weighted average of the absolute difference between average confidence and accuracy across all bins, weighted by the number of samples in each bin. A lower ECE indicates better calibration (perfect calibration yields ECE=0). Frameworks like UQ4CT aim to minimize ECE during fine-tuning. (Posocco & Bonnefoy, 2021)
  • ECE under Distribution Shift: This involves calculating ECE specifically on OOD datasets or under simulated distribution shifts to assess how calibration holds up under stress. Research suggests that factors like OOD prompts versus OOD responses can impact the calibration of components like reward models differently (LeVine et al., 2024; Huang et al., 2024b).
@@
    • Brier Score: Measures the mean squared difference between predicted probabilities and actual outcomes (0 or 1). Its other variants include weighted Brier score (Zhu et al., 2024).

### v2

@@ 4.4 Post-processing techniques
A fundamental principle of post-processing is to treat the LLMs output with a degree of skepticism, applying verification steps similar to how one might validate user input. Common strategies include safety classifiers, where smaller models or rule-based systems screen outputs for harmful or biased content, blocking or flagging problematic responses (kumar2025llmposttrainingdeepdive; liu2023promptingframeworkslargelanguage; zhou2024robustpromptoptimizationdefending). For example, the Erase-and-Check method iteratively tests modified versions of suspicious prompts to detect adversarial inputs, leveraging pre-generation filtering to block harmful content (kumar2025certifyingllmsafetyadversarial; phute2024llmselfdefenseself). Perplexity filtering identifies nonsensical outputs by measuring how natural generated text appears to a language model, flagging high-perplexity results as potential errors or attacks (weng2023attack).
@@
For structured tasks (e.g., code or json generation), format validation ensures outputs adhere to predefined schemas or pass external checks like code compilers, preventing errors or injection vulnerabilities (rebedea2023nemoguardrailstoolkitcontrollable; guardrailsai2025). Similarly, output sanitization removes sensitive information (e.g., personal data) leaked during generation, mitigating privacy risks (wang2025surveyresponsiblellmsinherent). Further, song2025assessing; maheshwari2025citefixenhancingragaccuracy implements RAG to verify outputs after initial response generation. After a language model produces an answer, the system automatically fetches supporting evidence from external databases or documents, using either the original user query or the generated response as search criteria. The model then cross-checks its own output against this retrieved evidence, creating a self-correction cycle that improves factual consistency. While methods like Erase-and-Check offer strong defences, their computational cost often requires simplified variants for practical use, highlighting the trade-off between robustness and efficiency. Together, these strategies underscore the importance of layered verification to balance safety, accuracy, and usability in real-world LLM deployments.
Table 3: Decision Matrix for Mitigation Strategy Selection
Threat Model (Source of Non-Robustness) Primary Dimension(s) Addressed Recommended Mitigation Strategies (by Pipeline Stage) Key Considerations / Trade-offs
Dataset Biases & Anomalies Fairness under Stress, Consistency Pre-processing: Data Filtering (Rule-based, Data-Juicer, LLM-as-cleaner), Data Augmentation (Contrastive pairs); Trade-off: Overly strict filtering can remove useful data.                    Cost: Manual data curation is high, but
In-processing: Alignment (Safe RLHF), Regularization (DBL) automated data curation is lower. Effectiveness: Crucial for ethical AI.
Data Poisoning / Backdoors ifdversarial Attacks, Safety, Reliability Pre-processing: Data Filtering (rigorous vetting); Trade-off: AT is expensive, attack-specific. Cost: High for AT.
In-processing: Adversarial Training (AT, DRO); Post-processing: Output Filtering (Safety Classifiers), LLM-as-a-Judge (detection) Effectiveness: Continuous monitoring needed. Scalability: Challenging for large models.
Sensitivity to Input Variations Prompt Robustness, Noise Robustness, Pre-processing: Data Augmentation (EDA, A3); Intra-processing: Robust Prompting Trade-off: Overly restrictive prompts stifle creativity.
(Prompt/Noise) Consistency (Self-denoising, Referenced Instruction Tracking), Inference-Time Adaptation Cost: Lower (inference-time) than retraining. Generalizability: Adapts to emerging risks.
RLHF Impact (Reward Hacking, Fairness under Stress, Consistency, In-processing: Alignment (RPO, Ensemble RMs, Consistency Alignment) Trade-off: “Alignment tax” (reduced general capabilities).
Bias Amplification, Capability Trade-offs) Task-Specific Robustness Cost: High computational resources. Scalability: Limited by data quality/diversity.
Alignment Tax OOD Generalization, Task-Specific Robustness In-processing: Alignment (RPO, Consistency Alignment, Model Averaging) Trade-off: Direct trade-off with general utility. Effectiveness: Aims to balance safety with performance.
Fine-tuning Limitations OOD Generalization, Noise Robustness, Pre-processing: Data Augmentation, Data Filtering; Trade-off: Risk of narrowing model’s broader capabilities.
(Overfitting, Noise Prompt Robustness In-processing: Regularization, AT; Cost: Depends on strategy.
Sensitivity, Instruction Brittleness) Intra-processing: Robust Prompting Effectiveness: Requires careful data curation.
RAG Issues Consistency, Factuality Intra-processing: Inference-Time Adaptation Trade-off: Retrieval overhead.
(Efficient inference engines); Cost: Adds latency.
Post-processing: RAG for Verification Effectiveness: Improves grounding, but bottlenecked by retriever quality.
@@
This approach uses one language model to evaluate the outputs of another, offering scalable and nuanced assessments beyond basic rule-based checks. This method is increasingly applied to enhance robustness by automating tasks like bias detection, safety compliance, and factual accuracy (positionshi2025judgingjudgessystematicstudy; gu2025surveyllmasajudge; kumar2025llmposttrainingdeepdive). For instance, judge LLMs can systematically analyze responses for fairness across demographic categories, validate reasoning steps in technical tasks (e.g., code vulnerability detection), or flag factual errors by cross-referencing trusted sources (cantini2025benchmarkingadversarialrobustnessbias). Advanced frameworks even deploy judge LLMs to autonomously generate adversarial prompts, test target models, and assess response reliability, guided by domain-specific constraints (li2025wantedknowllmbasedvulnerability; beyer2025llmsafetyevaluationslackrobustness).
However, reliance on LLM judges introduces challenges. First, their evaluations can be inconsistent due to prompt sensitivity, inherent biases, or a lack of grounding in objective criteria (gu2025surveyllmasajudge). Second, judge models themselves are prone to manipulation; studies show adversarial phrases or stylistic cues (e.g., phrases like “I think”) can skew their assessments (raina-etal-2024-llm; Lee2024AreLR). Finally, the lack of standardized evaluation protocols and risks like “preference leakage”, where judges favour outputs from models they were trained on, underscores the need for multi-judge ensembles, human oversight, and rigorous benchmarking to ensure reliability (beyer2025llmsafetyevaluationslackrobustness). While promising, LLM-as-a-Judge systems require careful design to balance automation with trustworthiness.
@@ 5.1 Metrics
Performance Degradation Accuracy Drop, F1 Drop (yuan2023revisiting) General Robustness Decrease in standard metrics on challenging vs clean data.
Attack Success Rate (ASR) (zhou2024robustpromptoptimizationdefending) Adversarial Resilience, Safety Percentage of successful adversarial attacks/jailbreaks
Compliance / Refusal Rate (varshney-etal-2024-art; zeng2025rootdefencestrategiesensuring) Safety, Alignment Frequency of generating harmful content or refusing prompts
OOD Performance OOD Accuracy/F1 score (yuan2023revisiting) OOD Generalization Task performance on designated out-of-distribution datasets.
ID-OOD Performance Gap (yuan2023revisiting) OOD Generalization Difference between in-distribution and out-of-distribution performance.
OOD Detection (AUROC, Energy Score) (lu2024recentadvancesooddetection) OOD Awareness Ability to distinguish in-distribution versus out-of-distribution inputs.
Consistency Semantic Consistency (yang2025enhancingsemanticconsistencylarge) Prompt Robustness, Reliability Stability of output meaning under input paraphrasing or slight modifications.
Consistency Rate (CR) (nalbandyan2025scoresystematicconsistencyrobustness) Prompt Robustness, Sampling Stability Proportion of identical or equivalent predictions across setups (e.g., prompts, seeds).
@@
(RC) (rahman2024defandefinitiveanswerdataset) paraphrased prompts.
@@
Consistency Score (saxena2024evaluatingconsistencyreasoningcapabilities) Reliability samples.
Calibration Expected Calibration Error (ECE) (posocco2021estimatingexpectedcalibrationerrors) Confidence Reliability Alignment between predicted confidence and actual accuracy
ECE under Distribution Shift (huang-etal-2024-calibrating; levine2024baselineanalysisrewardmodels) OOD Generalization ECE measured specifically on OOD data
Fairness under Stress Bias Metrics (e.g., FLEX) (jung2025flexbenchmarkevaluatingrobustness) Fairness Robustness Evaluation of fairness when models are exposed to bias-inducing prompts.
Task-Specific Code Robustness Checks (unknowntask) Code Generation Robustness Checks for robustness in generated code, like security or correctness features.
Reasoning Accuracy (e.g., ReClor-plus) (bao2025assessingenhancingrobustnesslarge) Reasoning Robustness Performance on reasoning tasks containing structure variations.
@@
(rahman2024defandefinitiveanswerdataset) Faithfulness knowledge source or context.
FEWL (wei2024measuringreducingllmhallucination) Factuality (Reference-Free) Weighted agreement among reference LLMs based on expertise.
SelfCheckGPT, MetaQA (wei2024measuringreducingllmhallucination) Factuality/Consistency (Reference-Free) (chen2024hallucinationdetectionrobustlydiscerning) Consistency checks across multiple generated samples or metamorphic prompt variations.
@@ 5.1.1 Performance Degradation Metrics
These metrics quantify the drop in performance when an LLM is evaluated on challenging data (adversarial or noisy) compared to its performance on clean, standard data. They provide a direct measure of how much a specific challenge impacts the model’s effectiveness.
  • Accuracy Drop / F1 Drop / Metric Drop: This metric calculates the difference in model performance between a baseline (clean or standard) dataset and a challenging version of that dataset (e.g., perturbed, adversarial, or noisy) yuan2023revisiting. This metric directly quantifies how much a given challenge degrades effectiveness when expressed in absolute terms or percentage points. A larger drop indicates lower robustness, while a smaller drop suggests resilience under stress.
  • Attack Success Rate (ASR): This metric is central to adversarial robustness evaluation. It measures the percentage of adversarial inputs that successfully cause the intended model failure, such as misclassification in NLU tasks, bypassing safety filters to generate harmful content, or successfully executing a jailbreak prompt (jung2025flexbenchmarkevaluatingrobustness). A higher ASR signifies lower robustness to the specific attack method used. It is a key metric for benchmarks like AdvGLUE++, JailbreakBench, and PromptBench (zhou2024robustpromptoptimizationdefending).
  • Compliance Ratio / Refusal Rate: These metrics are particularly relevant for evaluating safety and alignment robustness, especially against jailbreaking attempts. It measures how often the LLM generates harmful or forbidden content when prompted with malicious instructions (varshney-etal-2024-art; zeng2025rootdefencestrategiesensuring). Conversely, the Refusal Rate measures how often the model refuses to answer. A high refusal rate for harmful prompts indicates good safety alignment, but a high refusal rate for benign prompts indicates over-defensiveness, a potential negative side-effect of safety training.
@@ 5.1.2 Out-of-Distribution (OOD) Performance Metrics
  • OOD Accuracy / F1 score: Standard performance metrics like Accuracy or F1-score calculated directly on designated OOD benchmark datasets, such as those included in the BOSS suite (e.g., Dynasent, ANLI, NewsQA) (yuan2023revisiting). This provides a direct measure of capability on unseen distributions.
  • ID-OOD Performance Gap: This metric quantifies the difference in performance between an ID dataset and a corresponding OOD dataset (yuan2023revisiting). It explicitly measures the drop in generalization capability due to the distribution shift. Studies using benchmarks like BOSS have shown that this gap can vary significantly and exhibit complex patterns (e.g., linear, piecewise linear, V-shaped) depending on factors like model scale, training steps, and task type, indicating that simply improving ID performance does not always guarantee better OOD performance.
  • OOD Detection Metrics (e.g., AUROC, FPR@TPR): These metrics evaluate the model’s ability to recognize that an input is OOD, often by thresholding a confidence score derived from model outputs (e.g., token probabilities) or internal representations (huang2024trustllmtrustworthinesslargelanguage). Common metrics include the Area Under the Receiver Operating Characteristic curve (AUROC) and the False Positive Rate (FPR) at a high True Positive Rate (TPR). While not directly measuring task performance robustness, effective OOD detection is crucial for building reliable systems that can identify and potentially abstain from making predictions on inputs they are likely to handle poorly (lu2024recentadvancesooddetection). Techniques like Energy Score, originally from classification, have been adapted for this purpose in LLMs and reward models (levine2024baselineanalysisrewardmodels).
@@ 5.1.3 Consistency Metrics
Consistency metrics evaluate the stability of LLM outputs. A consistent model should produce similar outputs for semantically identical inputs and exhibit stable behaviour under non-deterministic sampling. Lack of consistency can signal unreliability, especially for tasks requiring factual accuracy (nalbandyan2025scoresystematicconsistencyrobustness).
  • Semantic Consistency: Assesses whether the meaning or core information conveyed in the LLM’s output remains the same when the input prompt is paraphrased or expressed differently, while preserving the original intent (nalbandyan2025scoresystematicconsistencyrobustness). Evaluation can be qualitative or quantitative, often using semantic similarity metrics based on embeddings (e.g., cosine similarity between output embeddings) (patwardhan2025automatedconsistencyanalysisllms).
  • Consistency Rate (CR): Proposed within the SCORE framework (nalbandyan2025scoresystematicconsistencyrobustness), CR quantifies the agreement among predictions for the same input across different evaluation settings (e.g., 10 different prompts, 5 different random seeds for sampling). It is calculated as the average proportion of agreeing pairs among all possible pairs of predictions for each input instance. For multiple-choice questions (MCQ), agreement means identical predicted letters; for tasks like MATH, it means symbolic equivalence checked via tools like sympy (nalbandyan2025scoresystematicconsistencyrobustness).
  • Response Consistency (RC): Used in the DefAn benchmark (bao2025assessingenhancingrobustnesslarge), this metric specifically measures consistency across multiple (e.g., 15) paraphrased versions of a single prompt. For each set of paraphrased prompts, it calculates the frequency of the most common factual claim generated in the responses. The overall RC is the average of these frequencies.
  • Multilingual Consistency: Evaluates whether the model’s performance or the semantic meaning of its output is preserved when prompts or inputs are translated into different languages (Qi_2023).
  • LLM-based Consistency Score: This approach uses another LLM as a judge (saxena2024evaluatingconsistencyreasoningcapabilities). It prompts the target LLM multiple times (n > 1) for the same input using a non-zero temperature (e.g., 1.0) to induce variability. The judge LLM then compares pairs of generated responses, assessing whether they contradict or support each other. The consistency score is the fraction of pairs deemed consistent.
@@ 5.1.4 Calibration Metrics
  • Expected Calibration Error (ECE): It works by dividing predictions into bins based on their confidence scores (e.g., 10 bins from 0.0-0.1, 0.1-0.2,…, 0.9-1.0). Within each bin, the average confidence and the actual accuracy (fraction of correct predictions) are calculated. ECE is the weighted average of the absolute difference between average confidence and accuracy across all bins, weighted by the number of samples in each bin. A lower ECE indicates better calibration (perfect calibration yields ECE=0). Frameworks like UQ4CT aim to minimize ECE during fine-tuning. (posocco2021estimatingexpectedcalibrationerrors)
  • ECE under Distribution Shift: This involves calculating ECE specifically on OOD datasets or under simulated distribution shifts to assess how calibration holds up under stress. Research suggests that factors like OOD prompts versus OOD responses can impact the calibration of components like reward models differently (levine2024baselineanalysisrewardmodels; huang-etal-2024-calibrating).
@@
    • Brier Score: Measures the mean squared difference between predicted probabilities and actual outcomes (0 or 1). Its other variants include weighted Brier score (zhu2024weightedbrierscore).